Every week i spend the same amount of money on post stamps. They cost 30cents currently. Next month cost will go down by 5cents. When this happens i will be able to buy 4 more every week. How much do i spend on stamps each week
Answer by CubeyThePenguin(3113)

x = amount of stamps bought weekly
cost = 30x
next month = 25(x + 4)

30x = 25(x + 4)
30x = 25x + 100
5x = 100
x = 20

cost = 30(20) = 600 cents = $6 a week